OpenStudy (blackops2_143):

Simplify. -4.5[2x - (0.5 - 1.4)] A. -9x + 0.9 B. -9x -0.9 C. -9x - 4.05 D. 9x - 4.05

OpenStudy (anonymous):

C. -9x - 4.05 −4.5(2x−(0.5−1.4)) =(−4.5)(2x+−0.5+1.4) =(−4.5)(2x)+(−4.5)(−0.5)+(−4.5)(1.4) =−9x+2.25−6.3 =−9x−4.05 Do you understand that?

rebeccaxhawaii (rebeccaxhawaii):

that is incorrect it would be D

OpenStudy (blackops2_143):

Thank you

OpenStudy (anonymous):

@rebeccaxhawaii Really? Oops...sorry! D:

rebeccaxhawaii (rebeccaxhawaii):

no im sorry i read the problem wrong you are correct good job

OpenStudy (anonymous):

Oh, that's okay. Thanks